# Nonlinear Multi-objective Reinforcement Learning with Provable Guarantees

### v2

@@ 4 Modeling Optimality
Note that an approximation algorithm for this discounted finite time horizon problem can also be used as an approximation algorithm for the discounted infinite time horizon problem. Informally, discounting by $\gamma$ with bounded maximum reward implies that the first $T\approx 1/(1-\gamma)$ steps dominate overall returns.
@@
$\bigg{\|}\sum_{k=1}^{\infty}\gamma^{k-1}\bm{R}(s_{k},a_{k})-\sum_{k=1}^{T}\gamma^{k-1}\bm{R}(s_{k},a_{k})\bigg{\|}$
$=$ $\bigg{\|}\sum_{k=T+1}^{\infty}\gamma^{k-1}\bm{R}(s_{k},a_{k})\bigg{\|}$
$=$ $\gamma^{T}\bigg{\|}\sum_{k=1}^{\infty}\gamma^{k-1}\bm{R}(s_{T+k},a_{T+k})\bigg{\|}$
@@
5 The RA-Value Iteration Algorithm
In this section, we describe a model-based algorithm called RA-Value Iteration that efficiently finds an approximately optimal policy. Note that the values of $V^{*}$ with $T$ timesteps remaining only depends on the those with $T-1$ timesteps remaining. This naturally induces a dynamic programming approach. A naive algorithm for learning a nonstationary policy would need to consider all possible prior trajectories for each decision point, leading to a runtime complexity that depends on the exponential term $|\mathcal{S}|^{T}$.
Instead, our algorithm is derived from the recursive definition of the optimal multi-objective value function $V^{*}$ in 7 justified by Lemma 6, parameterized by the accumulated discounted reward vector $\mathbf{R_{acc}}$ instead of the prior history. Note that even if rewards were integer, $\mathbf{R_{acc}}$ would not be due to discounting. We must therefore introduce a discretization method to our dynamic programming algorithm that treats “nearby” accumulated reward vectors as equivalent. It is parameterized by some $\alpha\in(0,1)$ where a smaller $\alpha$ leads to a finer discretization but increases the runtime.
Definition 9.
In other words, $f$ maps any $d$-dimension vector to its nearest $\alpha$ multiples. We are now prepared to define the algorithm.
Algorithm 1 The RA-Value Iteration Algorithm
1:Parameters: Discretization precision $\alpha\in(0,1)$, Discount factor $\gamma\in[0,1)$, Reward dimension $d$, finite time horizon $T$, uniformly continuous welfare function $W$, discretization function $f$.
2:Require: Normalize $\mathbf{R}(s,a)\in[0,1]^{d}$ for all $s\in\mathcal{S},a\in\mathcal{A}$. Initialize $V(s,\mathbf{R_{acc}},0)=W(\mathbf{R_{acc}})$ for all $\mathbf{R_{acc}}\in\{0,\mathbf{\alpha},2\mathbf{\alpha},\dots,\lfloor T/\alpha\rfloor\cdot\alpha\}^{d}$, $s\in\mathcal{S}$.
3:for t = 1 to $T$ do
4:     for $\mathbf{R_{acc}}\in\{0,\mathbf{\alpha},2\mathbf{\alpha},\dots,\lfloor(T-t)/\alpha\rfloor\cdot\alpha\}^{d}$ do
5:         for all $s\in\mathcal{S}$ do
6:
$V(s,\mathbf{R_{acc}},t)$ $\leftarrow\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times$
$\quad V(s^{\prime},f_{\alpha}(\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a)),t-1)$
7:
$\pi(s,\mathbf{R_{acc}},t)$ $\leftarrow\operatorname*{arg\,max}_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times$
$\quad V(s^{\prime},f_{\alpha}(\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a)),t-1)$
8:         end for
9:     end for
10:end for
The runtime of the RA-Value Iteration algorithm is $O\left(|\mathcal{S}||\mathcal{A}|(T/\alpha)^{d}\right)$ which is polynomial time for constant dimensionality $d$; recall that $d$ is the number of objectives, possibly just 1.
We now prove a worst-case approximation guarantee for the RA-Value Iteration algorithm: given a smooth welfare function $W$, the algorithm computes a non-stationary policy with approximately optimal expected welfare. We begin by proving an important structural property: the optimal value function inherits smoothness of the welfare function, i.e. the optimal value function will be smooth if the welfare function is smooth.
Lemma 10 (Uniform continuity of multi-objective value function).
[…]
@@ 5 The RA-Value Iteration Algorithm
We now prove the approximation guarantee of RA-Value Iteration, that it achieves an additive error that scales with the smoothness of $W$ and the number of time steps remaining.
Lemma 11 (Approximation Error of RA-Value Iteration).
@@
$\forall s\in\mathcal{S},\mathbf{R_{acc}}\in{\mathbb{R}}^{d},t\in\{0,1,\dots,T\}$, where $V(s,\mathbf{R_{acc}},t)$ is the value function calculated by Algorithm 1.
@@
We prove this additive error bound by induction on the number of timesteps $t$ remaining in the task.
[…]
$=V^{*}(s,\mathbf{R_{acc}},0)\quad\forall s,\mathbf{R_{acc}}$
Inductive hypothesis: for $t=0,\dots,k-1$,
$V(s,\mathbf{R_{acc}},t)\geq V^{*}(s,\mathbf{R_{acc}},t)-t\epsilon$
Then consider with $k$ steps remaining.
$V(s,\mathbf{R_{acc}},k)$
$=$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)V(s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a)),k-1)$
$\geq$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times$
$\bigg{(}V^{*}(s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a)),k-1)-(k-1)\epsilon\bigg{)}.$
where the last line follows from the inductive hypothesis. Since $W$ is uniformly continuous by Lemma 10, there exists $\delta_{\epsilon}$ such that
$\bigg{|}V^{*}\bigg{(}s^{\prime},f_{\alpha}(\mathbf{R_{acc}}$ $+\gamma^{T-k}\mathbf{R}(s,a)),k-1\bigg{)}$
$-V^{*}\bigg{(}s^{\prime},\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a),k-1\bigg{)}\bigg{|}<\epsilon$
@@
$\bigg{\|}f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a))-(\mathbf{R_{acc}}$ $+\gamma^{T-k}\mathbf{R}(s,a))\bigg{\|}$
$<(\delta_{\epsilon}/d)\cdot d<\delta_{\epsilon}.$
Then
$V(s,\mathbf{R_{acc}},k)$
$\geq$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)V^{*}\bigg{(}s^{\prime},\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a),k-1\bigg{)}$
$\quad-\epsilon-(k-1)\epsilon$

### v3

@@ 4 Modeling Optimality
Note that an approximation algorithm for this discounted finite time horizon problem can also be used as an approximation algorithm for the discounted infinite time horizon problem. Informally, discounting by $\gamma$ with bounded maximum reward implies that the first $T\approx 1/(1-\gamma)$ steps dominate overall returns. We defer the precise formulation of the lower bound on the horizon time and its proof in the appendix.
5 Computing Optimal Policies
Our overall algorithm is Reward-Aware Explore or Exploit (or RAEE for short) to compute an approximately optimal policy, inspired by the classical E3 algorithm [16]. At a high level, the algorithm explores to learn a model of the environment, periodically pausing to recompute an approximately optimal policy on subset of the environment that has been thoroughly explored. We call this optimization subroutine Reward-Aware Value Iteration (or RAVI for short). In both cases, reward-aware refers to the fact that the algorithms compute non-stationary policies in the sense that optimal behavior depends on currently accumulated vector-valued reward, in addition to the current state.
Of the two, the model-based optimization subroutine RAVI is the more significant. The integrated model-learning algorithm RAEE largely follows from prior work, given access to the RAVI subroutine. For this reason, we focus in this section on RAVI, and defer a more complete discussion and analysis of RAEE to Section 7 and the appendix.
A naive algorithm for computing a nonstationary policy would need to consider all possible prior trajectories for each decision point, leading to a runtime complexity containing the term $|\mathcal{S}|^{T}$, exponential in the size of the state space and time horizon. Instead, for a smooth welfare function on a constant number of objectives, our algorithm will avoid any exponential dependence on $|\mathcal{S}|$.
The algorithm is derived from the recursive definition of the optimal multi-objective value function $V^{*}$ in Definition 6, justified by Lemma 1, parameterized by the accumulated discounted reward vector $\mathbf{R_{acc}}$ instead of the prior history. Note that even if rewards were integer, $\mathbf{R_{acc}}$ might not be due to discounting. We must therefore introduce a discretization that maps accumulated reward vectors to points on a lattice, parameterized by some $\alpha\in(0,1)$ where a smaller $\alpha$ leads to a finer discretization but increases the runtime.
Definition 7.
In other words, $f$ maps any $d$-dimension vector to its nearest $\alpha$ multiples. We are now prepared to describe the algorithm, which at a high level, uses computes the dynamic program of approximately welfare-optimal value functions conditioned on discretized accumulated reward vectors.
Algorithm 1 Reward-Aware Value Iteration (RAVI)
1:Parameters: Discretization precision $\alpha\in(0,1)$, Discount factor $\gamma\in[0,1)$, Reward dimension $d$, finite time horizon $T$, welfare function $W$, discretization function $f_{\alpha}$.
2:Require: Normalize $\mathbf{R}(s,a)\in[0,1]^{d}$ for all $s\in\mathcal{S},a\in\mathcal{A}$.
3:Initialize: $V(s,\mathbf{R_{acc}},0)=W(\mathbf{R_{acc}})$ for all $\mathbf{R_{acc}}\in\{0,\mathbf{\alpha},2\mathbf{\alpha},\dots,\lceil T/\alpha\rceil\cdot\alpha\}^{d}$, $s\in\mathcal{S}$.
4:for t = 1 to $T$ do
5:     for $\mathbf{R_{acc}}\in\{0,\mathbf{\alpha},2\mathbf{\alpha},\dots,\lceil(T-t)/\alpha\rceil\cdot\alpha\}^{d}$ do
6:         for all $s\in\mathcal{S}$ do
7:
$V(s,\mathbf{R_{acc}},t)\leftarrow\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times V(s^{\prime},f_{\alpha}(\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a)),t-1)$
8:
$\pi(s,\mathbf{R_{acc}},t)\leftarrow\operatorname*{arg\,max}_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times V(s^{\prime},f_{\alpha}(\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a)),t-1)$
The asymptotic runtime complexity is $O\left(|\mathcal{S}|^{2}|\mathcal{A}|(T/\alpha)^{d}\right)$. However, observe that the resulting algorithm is extremely parallelizable: Given solutions to subproblems at $t-1$, all subproblems for $t$ can in principle be computed in parallel. .
The parallel computing implementation of the RAVI algorithm leverages the computational power of GPUs to handle the extensive calculations involved in multi-objective value iteration. Each thread computes the value updates and policy decisions for a specific state and accumulated reward combination, which allows for massive parallelism.
In practice, we observed an empirical speedup of approximately 560 times on a single NVIDIA A100 compared to the CPU implementation on an AMD Ryzen 9 7950x 16-core processor for our experimental settings. This drastic improvement in runtime efficiency makes it feasible to run RAVI on larger and more complex environments, where the computational demands would otherwise be prohibitive.
It remains to see how small $\alpha$ needs to be, which will dictate the final runtime complexity. We first analyze the correctness of the algorithm and then return to the setting of $\alpha$ for a given smoothness of $W$ to conclude the runtime analysis.
We begin analyzing the approximation of RAVI with an important structural property: the optimal value function is smooth if the welfare function is smooth. The proof is included in the appendix.
Lemma 2 (Uniform continuity of multi-objective value function).
$\bigg{|}V^{*}(s,\mathbf{R_{1}},t)-V^{*}(s,\mathbf{R_{2}},t)\bigg{|}<\epsilon\quad\text{if}\quad\|\mathbf{R_{1}}-\mathbf{R_{2}}\|<\delta_{\epsilon}.$
We now present the approximation guarantee of RAVI, that it achieves an additive error that scales with the smoothness of $W$ and the number of time steps remaining. The proof of this lemma is deferred to the appendix.
Lemma 3 (Approximation Error of RAVI).
$V(s,\mathbf{R_{acc}},t)\geq V^{*}(s,\mathbf{R_{acc}},t)-t\epsilon$
$\forall s\in\mathcal{S},\mathbf{R_{acc}}\in\mathbb{R}^{d},t\in\{0,1,\dots,T\}$, where $V(s,\mathbf{R_{acc}},t)$ is computed by Algorithm 1.
While we can use any setting of $\alpha$ empirically, this shows that for an approximation guarantee we should set the discretization parameter to $\alpha=\delta_{T\epsilon}/d$ where $\delta_{T\epsilon}$ from the proof of Lemma 3 is sufficient to drive its bound to $\epsilon$, that is, it should be $\alpha=\frac{\delta_{\epsilon}}{T\cdot d}.$
We thus arrive at the ultimate statement of the approximation and runtime of RAVI. The proof follows directly from Lemma 3 and the setting of $\alpha$.
Theorem 1 (Optimality Guarantee of RAVI).
For a given $\epsilon$ and welfare function $W$ that is $\delta_{\epsilon}$ uniformly continuous, RAVI with $\alpha=\frac{\delta_{\epsilon}}{T\cdot d}$ computes a policy $\hat{\pi}$, such that $V^{\hat{\pi}}(s,0,T)\geq V^{*}(s,0,T)-\epsilon$ in $O\left(|\mathcal{S}|^{2}|\mathcal{A}|(d\cdot T^{2}/\delta_{\epsilon})^{d}\right)$ time.
A concrete example of a particular welfare function, the setting of relevant parameters, and a derivation of a simplified runtime statement may help to clarify Theorem 1. Consider the smoothed proportional fairness objective: $W_{\text{SPF}}(\mathbf{x})=\sum_{i=1}^{d}\ln(x_{i}+1)$, a smoothed log-transform of the Nash Social Welfare (or geometric mean) with better numerical stability.
Uniform continuity is a weaker modeling assumption than $L$-Lipshitz continuity for constant $L$. Note that for an $L$-Lipshitz function, the correct value of $\delta_{\epsilon}$ is just $\epsilon/L$, where $\epsilon$ is the desired approximation factor of the algorithm. By taking the gradient, we can see that the $W_{SPF}(\mathbf{x})$ is $d$-Lipshitz for nonnegative $\mathbf{x}$, where $d$ is the dimension of vector reward, so $\delta_{\epsilon}=\epsilon/d$. In practice, as we will see later, this is actually pessimistic: The objective is significantly smoother farther away from the origin, and a discretization that grows increasingly coarse farther from the origin can still efficiently approximate the objective. Nevertheless, Theorem 1 on this example would give us a runtime of
$O\left(|\mathcal{S}|^{2}|\mathcal{A}|(T/\alpha_{T\epsilon})^{d}\right)=O\left(|\mathcal{S}||\mathcal{A}|\left(\frac{T^{2}\cdot d^{2}}{\epsilon}\right)^{d}\right).$
To further simplify, if one takes the number of actions $|\mathcal{A}|$, discount factor $\gamma$, and dimension $d$ to be small constants, then the asymptotic dependence of the runtime (in our running example) is
$O\left(|\mathcal{S}|^{2}\left(\frac{1}{\epsilon}\log^{2}(1/\epsilon)\right)^{d}\right).$
By contrast, a brute force algorithm that conditions on all previous trajectories, even with these assumptions must scale no better than $O(|\mathcal{S}|^{T})$.
6 Experiments and Empirical Observations
We test RAVI on two distinct interpretations of multi-objective reinforcement learning: A. the fairness interpretation where the agent tries to maximize rewards across all dimensions. B. the objective interpretation, where the agent tries to maximize one while minimizing the other. In both scenarios, we show that RAVI can discover more optimal policies than other baselines for nonlinear multi-objective optimization. This holds even for coarser settings of $\alpha$ in the algorithm than would be necessary for strong theoretical worst-case approximation guarantees.
6.1 Taxi
We use the taxi multi-objective simulation environment considered in [13] for testing nonlinear ESR maximization. In this grid-world environment (Figure 3), the agent is a taxi driver whose goal is to deliver passengers. There are multiple queues of passengers, each having a given pickup and drop-off location. Each queue is a different objective or dimension of the reward. The taxi can only take one passenger at a time and wants to balance trips for passengers in the different queues.
We adopt the Nash social welfare function $W_{\text{Nash}}(\mathbf{r})=\left(\prod_{i}r_{i}\right)^{1/d}$ and egalitarian welfare $W_{\text{Egalitarian}}(\mathbf{r})=\min\{r_{i}\}_{i}$ for fairness considerations. To ensure numerical stability, Nash Welfare is optimized with its smoothed log-transform $W_{\text{SPF}}(\mathbf{r},\lambda)=\sum_{i=1}^{d}\ln(r_{i}+\lambda)$ for $\lambda=10^{-7}$. During evaluations, we still use $W_{\text{Nash}}$.
We consider two intuitive baseline approaches for optimization in a resource allocation task like this one. A. Linearly Scalarized Policy [23]: A relatively straightforward technique for multi-objective RL optimization is to apply the linear combination on the Q-values for each objective. Given weights $\mathbf{w}\in\mathbb{R}^{d}$, $\sum_{i}^{d}w_{i}=1$, the scalarized objective is $SQ(s,a)=\mathbf{w}^{\top}\mathbf{Q}(s,a)$, where $Q(s,a)_{i}$ is the Q-value for the $i^{\text{th}}$ objective. B. Mixture Policy [22]: this baseline works by combining multiple Pareto optimal base policies into a single policy. The mixing is achieved by adopting a base policy for $I$ steps and then switching to the next.
All the hyperparameters, including the interval for mixture policy and the weights for linear scalarization, are tuned via grid search.
We run RAVI along with the baselines on 20 random initializations of a $10\times 10$ taxi environment with three passenger spawn locations and destinations ($d=3$), $T=100$ without reward discount ($\gamma=1$), and compare their scalarized total rewards at convergence. As shown in 1, we observe that RAVI is able to compute better performing policies across both nonlinear objectives compared to both baselines.
6.2 Scavenger Hunt
Inspired by the Resource Gathering environment [8], the scavenger hunt environment is a grid-world simulation where the agent must collect resources while avoiding enemies scattered across the grid. The state representation includes the agent’s position and the status of the resources (collected or not). The reward function is vector-valued, where the first component indicates the number of resources collected, and the second component indicates the damage taken by encountering enemies.
Figure 2: Illustration of the axi environment. The taxi is represented by the yellow circle and passenger spawn and delivery locations are represented by colored cells. Note that the scale of the actual experiments are much larger.
Figure 3: Illustration of the scavenger hunt environment. The agent is represented by the blue cell, the green cells are resources, and the red cells are enemies.
Nash Welfare Egalitarian Welfare RAVI 5.06$\pm$0.13 4.35$\pm$0.49 Mixture Policy 3.99$\pm$0.43 2.86$\pm$0.33 Linearly Scalarized Policy 0.02$\pm$0.26 0.03$\pm$0.14
Table 1: Comparison with baselines in terms of ESR for the taxi environment.
Cobb-Douglas RD Threshold RAVI 1.63$\pm$0.20 3.12$\pm$0.48 Welfare Q-learning 1.044$\pm$0.20 1.33$\pm$0.54 Linearly Scalarized Policy 1.02$\pm$0.32 1.68$\pm$0.50
Table 2: Comparison with baselines in terms of ESR for the scavenger hunt environment.
Given the differing nature of the task from the taxi environment, where the objective is to maximize one reward component while minimizing the other, we adopt two specific scalarization functions for the scavenger hunt environment: A. Resource-Damage (RD) Threshold Scalarization: $RD_{\text{threshold}}(R,D)=R-\max(0,(D-\text{threshold})^{3})$, for the number of resources collected $R$ and the damage taken $D$. The threshold parameter, i.e. the budget, determines the point at which the penalty on the damage taken starts to apply. The RD threshold scalarization encodes the concept of a budget into the optimization process. It is a constrained optimization approach with a soft penalty on the damage taken. B. Cobb-Douglas Scalarization: $CD(R,D)=R^{\alpha}\left(\frac{1}{D+1}\right)^{\beta}$. The Cobb-Douglas scalarization function is inspired by economic theory. It balances the trade-off between the number of resources collected and the damage taken. The parameters $\alpha$ and $\beta$ represent the weights assigned to each component, with $\alpha+\beta=1$. Note that when $\alpha$ and $\beta$ are set to 0.5, this scalarization function reduces $W_{\text{Nash}}(R,1/(D+1))$.
Due to the differing nature of the scavenger hunt task, a mixture policy does poorly — sometimes it ignores the resources and just tries not to hit an enemy, other times it ignores the enemies and just tries to get resources. We therefore compare the performance of RAVI against a linearly scalarized policy and Welfare Q-learning [13]: by using nonlinear updates and a non-stationary policy, this algorithm adapts the Q-learning algorithm to optimize ESR for a potentially nonlinear welfare function.
Similarly, we run both RAVI and Welfare Q-learning on 20 random initializations of a $15\times 15$ scavenger hunt environment with six resources and 40% of the cells randomly populated with enemies, $T=20$ without reward discount ($\gamma=1$), and compare their scalarized total rewards at convergence. As shown in table 2, we again observe that RAVI is able to compute better performing policies across both nonlinear objectives compared to Welfare Q-learning.
Discretization of Accumulated Rewards.
We include an ablation study on the setting of the discretization interval length $\alpha$ in the appendix, in wihch we test values of $\alpha$ are substantially greater (thus making the algorithm more runtime efficient) than what would be necessary for worst-case theoretical guarantees from the previous analysis. Note that the empirical performance of RAVI is substantially better than is guaranteed by the previous theoretical analysis, for which substantially smaller $\alpha$ values would be necessary for worst-case guarantees.
7 Removing Knowledge of the Model
We have shown that RAVI can efficiently find an approximately optimal policy given access to a model of the environment. In this section we observe that the model can be jointly learned by extending the classical E3 algorithm [16] to the nonlinear multi-objective case by lifting the exploration algorithm to the multi-objective setting and using RAVI for the exploitation subroutine. We call the resulting combined algorithm Reward-Aware-Explore or Exploit (or RAEE for short). We briefly explain the high level ideas and state the main result here and defer further discussion to the appendix due to space constraints.
The algorithm consists of two stages, exploration and exploitation. Repeat the following stages, described here at a high level and in the appendix in more detail.
  • Explore. At a given state, choose the least experienced action and record the reward and transition. Continue in this fashion until reaching a known state, where we say a state is known if it has been visited sufficiently many times for us to have precise local statistics about the reward and transition functions at that state.
  • Exploit. Run RAVI from the current known state $s$ on the induced model comprising the known states and with a single absorbing state representing all unknown states. If the welfare obtained by this policy is within the desired error bound of $V^{*}(s,\mathbf{0},T)$ then we are done. Otherwise, compute a policy to reach an unknown state as rapidly as possible and resume exploring.
Theorem 2 (RAEE).
Let $V^{*}(s,0,T)$ denote the value function for the policy with the optimal expected welfare in the MOMDP $M$ starting at state $s$, with $\mathbf{0}\in\mathbb{R}^{d}$ accumulated reward and $T$ timesteps remaining. Then for a uniformly continuous welfare function $W$, there exists an algorithm $A$, taking inputs $\epsilon$, $\beta$, $\mathcal{S}$, $\mathcal{A}$, and $V^{*}(s,\mathbf{0},T)$, such that the total number of actions and computation time taken by $A$ is polynomial in $1/\epsilon$, $1/\beta$, $|\mathcal{S}|$, $|\mathcal{A}|$, the horizon time $T=1/(1-\gamma)$ and exponential in the number of objectives $d$, and with probability at least $1-\beta$, $A$ will halt in a state $s$, and output a policy $\hat{\pi}$, such that $V^{\hat{\pi}}_{M}(s,0,T)\geq V^{*}(s,0,T)-\epsilon$.
We provide additional details comparing our analysis with that of [16] in the appendix.
8 Conclusion and Future Work
Nonlinear prefernces in reinforcement learning are important, as they can encode fairness as the nonlinear balancing of priorities across multiple objectives or risk attitudes with respect to even a single objective. Stationary policies are not necessarily optimal for such objectives. We derived an extended form of Bellman optimality to characterize the structure of optimal policies conditional on accumulated reward. We introduced the RAVI and RAEE algorithms for efficiently computing an approximately optimal policy. Our primary contribution is a theoretical analysis, but we have also supplemented with an experimental demonstration.
Our results introduce several natural directions for future work. On the technical side, one could try to handle the case of stochastic reward functions or a large number of objectives. Another direction would involve incorporating function approximation with deep neural networks into the algorithms to enable scaling to even larger state spaces and generalizing between experiences. Our theory suggests that it may be possible to greatly enrich the space of possible policies that can be efficiently achieved in these settings by conditioning function approximators on accumulated reward rather than necessarily considering sequence models over arbitrary past trajectories; we see this as the most exciting next step for applications.
Appendix
Appendix A Proofs of Technical Lemmas
In this section, we provide proofs of the technical lemmas in Section 4 and Section 5.
@@ Appendix A Proofs of Technical Lemmas
$\bigg{\|}\sum_{k=1}^{\infty}\gamma^{k-1}\mathbf{R}(s_{k},a_{k})-\sum_{k=1}^{T}\gamma^{k-1}\mathbf{R}(s_{k},a_{k})\bigg{\|}$
$=$ $\bigg{\|}\sum_{k=T+1}^{\infty}\gamma^{k-1}\mathbf{R}(s_{k},a_{k})\bigg{\|}$
$=$ $\gamma^{T}\bigg{\|}\sum_{k=1}^{\infty}\gamma^{k-1}\mathbf{R}(s_{T+k},a_{T+k})\bigg{\|}$
@@
Lemma 1. Let $\mathcal{V}(s,\mathbf{R}(\tau),0)=W(\mathbf{R}(\tau))$ for all states $s$ and trajectories $\tau$. For every state $s$, history $\tau$, and $t>0$ time steps remaining, let
Then $V^{*}(s,\tau,t)=\mathcal{V}(s,\mathbf{R}(\tau),t)$.
Proof.
We proceed by induction on $t$. In the base case of $t=0$, we have simply that $\mathcal{V}(s,\mathbf{R}(\tau),0)=W(\mathbf{R}(\tau))$ by the definition of $\mathcal{V}$. But when $t=0$, the trajectory has ended, so any $\tau^{\prime}$ in Definition 5 will be the empty trajectory $\tau_{0;0}$. So
Suppose the equality holds for all states $s$ and histories $\tau$ for up to $t-1$ time steps remaining, i.e.
Then with $t$ steps remaining:
$V^{*}(s,\tau,t)=\max_{\pi}\sum_{\tau^{\prime}}Pr^{\pi}(\tau^{\prime})W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(\tau^{\prime}))$
where the sum is taken over all length-$t$ trajectories $\tau^{\prime}$ beginning at $s$, and $Pr^{\pi}(\tau^{\prime})$ is the probability that $\tau^{\prime}$ is traversed under policy $\pi$. Note that we can further decompose $\tau^{\prime}$ into $\tau^{\prime}=(s,a)\oplus\tau^{\prime\prime}$ where $(s,a)$ denotes the state-action pair corresponding to the current timestep, $\oplus$ concatenates, and $\tau^{\prime\prime}$ denotes a trajectory of length $t-1$ beginning at some state $s^{\prime}$ transitioned from state $s$ via taking action $a$. Then we can rewrite $V^{*}$ as
$V^{*}(s,\tau,t)=\max_{\pi}\sum_{a}\pi(a\mid s,\tau,t)~{}\mathbb{E}_{s^{\prime}}\bigg{[}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a)+\gamma^{T-t+1}\mathbf{R}(\tau^{\prime\prime}))\bigg{]}$
$\mathbb{E}_{s^{\prime}}\bigg{[}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a)+\gamma^{T-t+1}\mathbf{R}(\tau^{\prime\prime}))\bigg{]}$
$V^{*}(s,\tau,t)$ $=\max_{a}\mathbb{E}_{s^{\prime}}\bigg{[}\max_{\pi}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a)+\gamma^{T-(t-1)}\mathbf{R}(\tau^{\prime\prime}))\bigg{]}$
$=\max_{a}\mathbb{E}_{s^{\prime}}\left[\mathcal{V}(s^{\prime},\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a),t-1)\right]$
$=\mathcal{V}(s,\mathbf{R}(\tau),t).$
Lemma 2. Let the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$ be uniformly continuous. Fix $s\in\mathcal{S}$ and $t\in\{0,1,\dots,T\}$, then for all $\epsilon>0$, there exists $\delta_{\epsilon}>0$ such that
$\bigg{|}V^{*}(s,\mathbf{R_{1}},t)-V^{*}(s,\mathbf{R_{2}},t)\bigg{|}<\epsilon\quad\text{if}\quad\|\mathbf{R_{1}}-\mathbf{R_{2}}\|<\delta_{\epsilon}.$
@@
Lemma 3 (Approximation Error of RAVI)
@@
$\forall s\in\mathcal{S},\mathbf{R_{acc}}\in\mathbb{R}^{d},t\in\{0,1,\dots,T\}$, where $V(s,\mathbf{R_{acc}},t)$ is computed by Algorithm 1.
@@
We prove this additive error bound by induction on the number of timesteps $t$ remaining in the task. Cleary $V(s,\mathbf{R_{acc}},0)=W(\mathbf{R_{acc}}),\forall s$ and $\mathbf{R_{acc}}$. Suppose that $V(s,\mathbf{R_{acc}},t)\geq V^{*}(s,\mathbf{R_{acc}},t)-t\epsilon$ for all $t=0,\dots,k-1$. Then consider with $k$ steps remaining.
$V(s,\mathbf{R_{acc}},k)=$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)V(s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a)),k-1)$
$\geq$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times\bigg{(}V^{*}(s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a)),k-1)-(k-1)\epsilon\bigg{)}.$
Since $W$ is uniformly continuous, by Lemma 2 there exists $\delta_{\epsilon}$ such that
$\bigg{|}V^{*}\bigg{(}s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a)),k-1\bigg{)}-V^{*}\bigg{(}s^{\prime},\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a),k-1\bigg{)}\bigg{|}<\epsilon$
@@
$\bigg{\|}f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a))-(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a))\bigg{\|}<(\delta_{\epsilon}/d)\cdot d=\delta_{\epsilon},$
and so
$V(s,\mathbf{R_{acc}},k)\geq$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)V^{*}\bigg{(}s^{\prime},\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a),k-1\bigg{)}-\epsilon-(k-1)\epsilon$